# labelsep = period

### v1

@@ 2.5 Lower bounds for linear systems of equations
The proof of the first result is immediate from the spectral sum lower bound (Theorem˜2.6) together with the trace estimation upper bound (Proposition˜2.2). The second result requires another (standard) ingredient, iterative refinement (Proposition˜D.1). See Section˜D.2 for the full proofs.
An interesting feature of this result is that the hard instance $\bm{A}$ is a symmetric, for which there is no distinction matvecs with $\bm{A}$ and $\bm{A}^{\smash{\intercal}}$. In particular, our results show that solving linear systems requires $\Omega(\kappa\log(\nicefrac{{1}}{{\varepsilon}}))$ matvecs even when $\bm{A}$ is symmetric. Of course, solving a linear system $\bm{A}\bm{x}=\bm{b}$ for a general matrix $\bm{A}$ can only be more difficult than the symmetric case, so the lower bound also applies against general matrices $\bm{A}$ with condition number $\kappa$ as well.
@@ 2.6 Inapproximability of the inverse function on a split interval
Imported Lemma 2.9 (Inapproximability of $\nicefrac{{1}}{{x}}$ on interval; Kraus et al. [24, Cor. 2.2] ).
@@ 3.2 Lower bounds for linear systems
By Yao’s minimax theorem [30, Sec. 2.2], we can assume without loss of generality that the algorithm is deterministic. Consider a one-sided matrix–vector algorithm performing $t$ matrix–vector products with transcript $\mathcal{T}$}, which we can assume without loss of generality are linearly independent. Instantiate the matrices $\bm{Q}$, $\bm{U}$, $\bm{W}$, and $\bm{V}$ from the hidden Haar theorem (Theorem˜3.1). The solution to $\bm{Q}\bm{x}=\mathbf{e}_{1}$ is
@@ 4 Conclusion
In this paper, we resolve the worst-case matrix-vector complexity of solving linear systems $\bm{A}\bm{x}=\bm{b}$ for one- and two-sided matrix–vector algorithms, up to small constant factors. Along the way, we also proved a lower bound on the complexity of estimating the inverse-trace $\operatorname{tr}(\bm{A}^{-1})$. Several natural questions remain to be explored in this space. Most directly, can further work on polynomial inapproximability enable similar matching upper and lower bounds for the $f(\bm{A})\bm{b}$ and $\operatorname{tr}(f(\bm{A}))$ problems for a much wider family of functions $f$?
@@ Acknowledgments
MD was supported in part by NSF CAREER Grant CCF-233865 and a Google ML and Systems Junior Faculty Award. ENE acknowledges support by the Miller Institute for Basic Research in Science, University of California Berkeley. RM would like to acknowledge the DARPA DIAL and DARPA AIQ programs for providing partial support of this work. This work was done in part while the authors were visiting the Simons Institute for the Theory of Computing. We thank Ainesh Bakshi, Chris Camaño, Cecilia Chen, Lin Lin, Michael Mahoney, Shyam Narayanan, and John Urschel for helpful conversations.
@@ A.1 Review: Total variation distance
Imported Lemma A.4 (Couplings lemma; Levin and Peres [25, Prop. 4.7]).
@@ A.3 Ingredient 2: From block Krylov to bilinear power sequences
Combining the ingredients to this point, we obtain the a general hardness result.
@@
By Yao’s minimax principle [30, Sec. 2.2], we can restrict our attention to deterministic algorithms. Consider any matrix–vector algorithm which can make $\ell$ matvecs, and let ${\mathcal{T}}=(\bm{A}\bm{x}_{i}\mathrel{\mathop{\ordinarycolon}}1\leq i\leq\ell)$ and $\tilde{}{\mathcal{T}}=(\tilde{}\bm{A}\tilde{}\bm{x}_{i}\mathrel{\mathop{\ordinarycolon}}1\leq i\leq\ell)$ denote transcripts of matvecs produced by the algorithm on $\bm{A}$ and $\tilde{}\bm{A}$, respectively. By the data processing inequality ˜A.1, the reduction to triangular block Krylov (Imported Theorem˜A.5), the observation that triangular block Krylov is a subset of a general block Krylov method, and the reduction to bilinear power sequences (Imported Lemma˜A.6), it holds that
@@ A.4 Ingredient 3: Constructing a finite set of eigenvalues
Imported Theorem A.8 (Inapproximability on a finite set; Rivlin [37, Thm. 1.14]).
@@ A.7 Ingredient 6: Total variation distance between Wishart matrices
Imported Lemma A.16 (Wishart vs. GOE; Rácz and Richey [38, Thm. 1.2]).
@@
Imported Lemma A.17 (TV between Gaussians; Devroye et al. [14, Thm. 1.3]).
@@ Appendix C Proof of the fine-grained lower bound
In this section, we establish Theorem˜1.3, which establishes a more fine-grained lower bound on the two-sided matrix–vector complexity of linear systems, as parametrized by a number $k$ of outlying singular values and a reduced condition number $\kappa_{k}=\sigma_{k+1}(\bm{A})/\sigma_{\mathrm{min}}(\bm{A})$. We use Theorem˜1.1 together with a reduction from Dereziński et al. [13, Lem. 34].
@@
Let $\delta>0$ be the universal constant in Imported Theorem˜C.1, and let $\bm{b}_{1},\ldots,\bm{b}_{k}$ be iid vectors, each with iid Rademacher entries. Then by standard analysis of the Girard–Hutchinson trace estimator [29, Lem. 2.1], setting $k={\mathcal{O}}(\log(1/\delta))={\mathcal{O}}(1)$ suffices to achieve
@@
Suppose there is an algorithm that attains the residual guarantee ˜C.1 while using at most at most $t$ matvecs. By taking the geoemtric median of ${\mathcal{O}}(\log(k/\delta))={\mathcal{O}}(1)$ independent executions of the algorithm, the success probability of the algorithm can be boosted to $1-\delta/2k$ at the cost of inflating the residual by a factor of at most $1.5$ [28, Thm. 3.1]. Applying this success-probability boosted procedure to each $\bm{b}_{i}$ (and noting that $\lVert\bm{b}_{i}\rVert_{2}=\sqrt{n}$), we obtain vectors $\tilde{}\bm{x}_{i}$ such that, for each $i$,
@@
Finally, combining with guarantee ˜C.2 and union bounding over each bound succeeding, we see that with proability at least $1-\delta$, it holds that
@@
The proof of Theorem˜1.3 is a corollary of the $\Omega(\kappa\log(1/\varepsilon))$ matvecs lower bound in Theorem˜1.1, the $\Omega(n)$ lower bound in ˜C.2, and the following result from the prior work:
@@
Fix any set of square matrices ${\mathcal{A}}\subseteq{\mathbb{R}}^{n\times n}$. Then the matrix–vector complexity for solving linear systems in $\mathcal{A}$}, denoted $T_{{\mathcal{A}}}(\varepsilon)$, is the smallest number $t$ such that there exist a two-sided matrix–vector algorithm which computes at most $t$ matrix–vector products and returns a vector $\tilde{}\bm{x}$ such that
@@
That is, the worst-case fine-grained matvec complexity of linear system solving is the sum of the matvec complexity of solving an arbitrary $k\times k$ linear system plus the matvec complexity of solving a $(n-k)\times(n-k)$ linear system with condition number $\kappa_{k}$. We can now prove Theorem˜1.3.
Proof of Theorem˜1.3.
@@
$T_{{\mathcal{F}}(n,k,\kappa_{k})}(\varepsilon)=\Omega(k^{3}+\kappa_{k}\log(1/\varepsilon)).\qed$
@@ Appendix D Deferred proofs
We let $\bm{b}\in{\mathbb{R}}^{n}$ be a uniformly random vector such that $\lVert\bm{b}\rVert_{2}=\sqrt{n}$. Then, by [19, Prop. 3.1] [^9] This article is in French. See, e.g., [18, Fact. 13.2] for a statement in English., we know that
@@
This type of result and its proof are standard [20, Sec. 2.1]. The idea is to use the oracle to produce an approximate solution $\tilde{}\bm{x}_{1}\approx\bm{A}^{-1}\bm{b}$. Then, we form the residual $\bm{r}_{1}=\bm{b}-\bm{A}\tilde{}\bm{x}_{1}$, use the oracle to produce an approximate solution $\tilde{}\bm{\delta}_{1}\approx\bm{A}^{-1}\bm{r}_{1}$, and form the corrected solution $\tilde{}\bm{x}_{2}=\tilde{}\bm{x}_{1}+\bm{\delta}_{1}$. Continuing this procedure for $r$ rounds produces a solution $\tilde{}\bm{x}_{r}$ satisfying the guarantee; we omit a detailed proof. Using this result, we prove Theorem˜2.8.
@@
Appendix E Additional upper and lower bounds
@@ E.1 $\ell_{2}$ error guarantees
The proof of Corollary˜E.1 is omitted; the proof is nearly the exact same as the proof of Theorem˜2.8 found in Section˜D.2. The key step is realizing the guarantee Corollary˜E.1 implies
@@
which was the same bound established in the proof of 2.8. Corollary˜E.1 shows that resolving a linear system to relative forwards error $\varepsilon$ also requires $\Omega(\kappa\log(\nicefrac{{1}}{{\varepsilon}}))$ matvecs. The best-known upper bounds for achieving an $\varepsilon$-small relative forward error requires at most ${\mathcal{O}}(\kappa\log(\kappa/\varepsilon))$ matvecs. The extra $\log(\kappa)$ factor comes from a comparison between the norms $\lVert\bm{A}^{-1}\bm{b}\rVert_{2}$ and $\lVert\bm{b}\rVert_{2}$. To sharpen the lower bound to match the upper bound, it would suffice to produce a lower bound of $\kappa\log(\kappa)$ matvecs being needed in the low-accuracy regime $\varepsilon\asymp 1$.
@@ E.2 Lower bound for SPD systems
Lastly, we show a $\Omega(\kappa\log(\nicefrac{{1}}{{\varepsilon}}))$ lower bound for SPD matrices. A $\Omega(\kappa\log(\nicefrac{{1}}{{\varepsilon}}))$ lower bound is known for deterministic algorithms [34, Sec. 7.2] and a $\Omega(\kappa/\operatorname{polylog}(\kappa))$ lower bound is known for randomized algorithms [5].
@@
Fix any $\delta,\eta\in(0,1)$ and suppose that $n\gtrsim\kappa^{(6+\delta)/\eta}$. Any algorithm using fewer than $\frac{1-\eta}{4}\sqrt{\kappa}\log n$ two-sided matvecs with a matrix $\bm{A}$ cannot return a vector $\tilde{}\bm{x}$ such that
@@
for the same class of problems requires $\Omega(\kappa)$ matvecs.
@@
Lower bounds for the forwards error in the $\ell_{2}$- and $\bm{A}$-norms are also possible, showing that $\Omega(\sqrt{\kappa}\log(1/\varepsilon))$ matvecs are needed in both cases.

### v2

@@ 2.5 Lower bounds for linear systems of equations
The proof of the first result is immediate from the spectral sum lower bound (Theorem˜2.6) together with the trace estimation upper bound (Proposition˜2.2). The second result requires another ingredient, iterative refinement (Proposition˜D.1). See Section˜D.2 for the full proofs.
An interesting feature of this result is that the hard instance $\bm{A}$ is symmetric, for which there is no distinction between matvecs with $\bm{A}$ and $\bm{A}^{\smash{\intercal}}$. In particular, our results show that solving linear systems requires $\Omega(\kappa\log(\nicefrac{{1}}{{\varepsilon}}))$ matvecs even when $\bm{A}$ is symmetric. Of course, solving a linear system $\bm{A}\bm{x}=\bm{b}$ for a general matrix $\bm{A}$ can only be more difficult than the symmetric case, so the lower bound also applies against general matrices $\bm{A}$ with condition number $\kappa$ as well.
@@ 2.6 Inapproximability of the inverse function on a split interval
Imported Lemma 2.9 (Inapproximability of $\nicefrac{{1}}{{x}}$ on interval; Kraus et al. [26, Cor. 2.2] ).
@@ 3.2 Lower bounds for linear systems
By Yao’s minimax theorem [33, Sec. 2.2], we can assume without loss of generality that the algorithm is deterministic. Consider a one-sided matrix–vector algorithm performing $t$ matrix–vector products with transcript $\mathcal{T}$}, which we can assume without loss of generality are linearly independent. Instantiate the matrices $\bm{Q}$, $\bm{U}$, $\bm{W}$, and $\bm{V}$ from the hidden Haar theorem (Theorem˜3.1). The solution to $\bm{Q}\bm{x}=\mathbf{e}_{1}$ is
@@ 4 Conclusion
In this paper, we resolve the worst-case matrix-vector complexity of solving linear systems $\bm{A}\bm{x}=\bm{b}$ for one- and two-sided matrix–vector algorithms, up to small constant factors. Along the way, we also prove a lower bound on the complexity of estimating the inverse-trace $\operatorname{tr}(\bm{A}^{-1})$. Several natural questions remain to be explored in this space. Most directly, can further work on polynomial inapproximability enable similar matching upper and lower bounds for the $f(\bm{A})\bm{b}$ and $\operatorname{tr}(f(\bm{A}))$ problems for a much wider family of functions $f$?
@@ Acknowledgments
MD was supported in part by NSF CAREER Grant CCF-233865 and a Google ML and Systems Junior Faculty Award. ENE acknowledges support by the Miller Institute for Basic Research in Science, University of California Berkeley. RAM would like to acknowledge the DARPA DIAL and DARPA AIQ programs for providing partial support of this work. This work was done in part while the authors were visiting the Simons Institute for the Theory of Computing. We thank Ainesh Bakshi, Chris Camaño, Cecilia Chen, Lin Lin, Michael Mahoney, Shyam Narayanan, and John Urschel for helpful conversations. Special thanks go to Ilse Ipsen for pointing us to the literature on information-based complexity.
@@ A.1 Review: Total variation distance
Imported Lemma A.4 (Couplings lemma; Levin and Peres [27, Prop. 4.7]).
@@ A.3 Ingredient 2: From block Krylov to bilinear power sequences
Combining the ingredients to this point, we obtain a general hardness result.
@@
By Yao’s minimax principle [33, Sec. 2.2], we can restrict our attention to deterministic algorithms. Consider any matrix–vector algorithm which can make $\ell$ matvecs, and let ${\mathcal{T}}=(\bm{A}\bm{x}_{i}\mathrel{\mathop{\ordinarycolon}}1\leq i\leq\ell)$ and $\tilde{}{\mathcal{T}}=(\tilde{}\bm{A}\tilde{}\bm{x}_{i}\mathrel{\mathop{\ordinarycolon}}1\leq i\leq\ell)$ denote transcripts of matvecs produced by the algorithm on $\bm{A}$ and $\tilde{}\bm{A}$, respectively. By the data processing inequality ˜A.1, the reduction to triangular block Krylov (Imported Theorem˜A.5), the observation that triangular block Krylov is a subset of a general block Krylov method, and the reduction to bilinear power sequences (Imported Lemma˜A.6), it holds that
@@ A.4 Ingredient 3: Constructing a finite set of eigenvalues
Imported Theorem A.8 (Inapproximability on a finite set; Rivlin [41, Thm. 1.14]).
@@ A.7 Ingredient 6: Total variation distance between Wishart matrices
Imported Lemma A.16 (Wishart vs. GOE; Rácz and Richey [42, Thm. 1.2]).
@@
Imported Lemma A.17 (TV between Gaussians; Devroye et al. [15, Thm. 1.3]).
@@ Appendix C Proof of the fine-grained lower bound
In this section, we establish Theorem˜1.4, which gives a more fine-grained lower bound on the two-sided matrix–vector complexity of linear systems, as parametrized by a number $k$ of outlying singular values and a reduced condition number $\kappa_{k}=\sigma_{k+1}(\bm{A})/\sigma_{\mathrm{min}}(\bm{A})$. We use Theorem˜1.1 together with a reduction from Dereziński et al. [13, Lem. 34].
@@
Let $\delta>0$ be the universal constant in Imported Theorem˜C.1, and let $\bm{b}_{1},\ldots,\bm{b}_{k}$ be iid vectors, each with iid Rademacher entries. Then by standard analysis of the Girard–Hutchinson trace estimator [32, Lem. 2.1], setting $k={\mathcal{O}}(\log(1/\delta))={\mathcal{O}}(1)$ suffices to achieve
@@
Suppose there is an algorithm that attains the residual guarantee ˜C.1 while using at most at most $t$ matvecs. By taking the geometric median of ${\mathcal{O}}(\log(k/\delta))={\mathcal{O}}(1)$ independent executions of the algorithm, the success probability of the algorithm can be boosted to $1-\delta/2k$ at the cost of inflating the residual by a factor of at most $1.5$ [31, Thm. 3.1]. Applying this success-probability boosted procedure to each $\bm{b}_{i}$ (and noting that $\lVert\bm{b}_{i}\rVert_{2}=\sqrt{n}$), we obtain vectors $\tilde{}\bm{x}_{i}$ such that, for each $i$,
@@
Finally, combining with guarantee ˜C.2 and union bounding over each bound succeeding, we see that with probability at least $1-\delta$, it holds that
@@
The proof of Theorem˜1.4 is a corollary of the $\Omega(\kappa\log(1/\varepsilon))$ matvecs lower bound in Theorem˜1.1, the $\Omega(n)$ lower bound in ˜C.2, and a result from prior work explained below.
@@
Fix any set of square matrices ${\mathcal{A}}\subseteq{\mathbb{R}}^{n\times n}$. Then the matrix–vector complexity for solving linear systems in $\mathcal{A}$}, denoted $T_{{\mathcal{A}}}(\varepsilon)$, is the smallest number $t$ such that there exists a two-sided matrix–vector algorithm which computes at most $t$ matrix–vector products and returns a vector $\tilde{}\bm{x}$ such that
@@
That is, the worst-case fine-grained matvec complexity of linear system solving is the sum of the matvec complexity of solving an arbitrary $k\times k$ linear system plus the matvec complexity of solving a $(n-k)\times(n-k)$ linear system with condition number $\kappa_{k}$. We can now prove Theorem˜1.4.
Proof of Theorem˜1.4.
@@
$T_{{\mathcal{F}}(n,k,\kappa_{k})}(\varepsilon)=\Omega(k+\kappa_{k}\log(1/\varepsilon)).\qed$
@@ Appendix D Deferred proofs
We let $\bm{b}\in{\mathbb{R}}^{n}$ be a uniformly random vector such that $\lVert\bm{b}\rVert_{2}=\sqrt{n}$. Then, by [21, Prop. 3.1] [^8] This article is in French. See, e.g., [20, Fact. 13.2] for a statement in English., we know that
@@
This type of result and its proof are standard [22, Sec. 2.1]. The idea is to use the oracle to produce an approximate solution $\tilde{}\bm{x}_{1}\approx\bm{A}^{-1}\bm{b}$. Then, we form the residual $\bm{r}_{1}=\bm{b}-\bm{A}\tilde{}\bm{x}_{1}$, use the oracle to produce an approximate solution $\tilde{}\bm{\delta}_{1}\approx\bm{A}^{-1}\bm{r}_{1}$, and form the corrected solution $\tilde{}\bm{x}_{2}=\tilde{}\bm{x}_{1}+\bm{\delta}_{1}$. Continuing this procedure for $r$ rounds produces a solution $\tilde{}\bm{x}_{r}$ satisfying the guarantee; we omit a detailed proof. Using this result, we prove Theorem˜2.8.
@@
Appendix E Additional lower bounds
@@ E.1 $\ell_{2}$ error guarantees
The proof of Corollary˜E.1 is omitted; the proof is nearly the same as the proof of Theorem˜2.8 found in Section˜D.2. The key step is realizing the guarantee Corollary˜E.1 implies
@@
which was the same bound established in the proof of 2.8. Corollary˜E.1 shows that resolving a linear system to relative forward error $\varepsilon$ also requires $\Omega(\kappa\log(\nicefrac{{1}}{{\varepsilon}}))$ matvecs. The best-known upper bounds for achieving an $\varepsilon$-small relative forward error require at most ${\mathcal{O}}(\kappa\log(\kappa/\varepsilon))$ matvecs. The extra $\log(\kappa)$ factor comes from a comparison between the norms $\lVert\bm{A}^{-1}\bm{b}\rVert_{2}$ and $\lVert\bm{b}\rVert_{2}$. To sharpen the lower bound to match the upper bound, it would suffice to produce a lower bound of $\kappa\log(\kappa)$ matvecs being needed in the low-accuracy regime $\varepsilon\asymp 1$.
@@ E.2 Lower bound for SPD systems
Lastly, we show an $\Omega(\sqrt{\kappa}\log(\nicefrac{{1}}{{\varepsilon}}))$ lower bound for SPD matrices. An $\Omega(\sqrt{\kappa}\log(\nicefrac{{1}}{{\varepsilon}}))$ lower bound is known for deterministic algorithms [38, Sec. 7.2] and an $\Omega(\sqrt{\kappa}/\operatorname{polylog}(\kappa))$ lower bound is known for randomized algorithms [5].
@@
Fix any $\delta,\eta\in(0,1)$ and suppose that $n\gtrsim\kappa^{(6+\delta)/\eta}$. Any randomized algorithm using fewer than $\frac{1-\eta}{4}\sqrt{\kappa}\log n$ two-sided matvecs with a matrix $\bm{A}$ cannot return a vector $\tilde{}\bm{x}$ such that
@@
for the same class of problems requires $\Omega(\sqrt{\kappa})$ matvecs.
@@
Lower bounds for the forward error in the $\ell_{2}$- and $\bm{A}$-norms are also possible, showing that $\Omega(\sqrt{\kappa}\log(1/\varepsilon))$ matvecs are needed in both cases.